# Domain-Invariant Representation Learning of Bird Sounds

### v5

@@ Abstract
Passive acoustic monitoring (PAM) is crucial for bioacoustic research, enabling non-invasive species tracking and biodiversity monitoring. Citizen science platforms like Xeno-Canto provide large annotated datasets from focal recordings, where the target species is intentionally recorded. However, PAM requires monitoring in passive soundscapes, creating a domain shift between focal and passive recordings, which challenges deep learning models trained on focal recordings. To address this, we leverage supervised contrastive learning to improve domain generalization in bird sound classification, enforcing domain invariance across same-class examples from different domains. We also propose ProtoCLR (Prototypical Contrastive Learning of Representations), which reduces the computational complexity of the SupCon loss by comparing examples to class prototypes instead of pairwise comparisons. Additionally, we present a new few-shot classification evaluation based on BIRB, a large-scale bird sound benchmark to evaluate bioacoustic pre-trained models. [^1] Models and code: https://github.com/ilyassmoummad/ProtoCLR
Index Terms:
Supervised Contrastive Learning, Domain Generalization, Few-shot Learning, Bioacoustics.
I Introduction
@@
However, a challenge arises from the difference between data collected on platforms like XC and PAM recordings. In citizen science platforms, recordings are typically focal—where the recorder is aimed directly at the species of interest. In contrast, PAM systems passively capture sounds within natural soundscapes, leading to recordings that contain a mix of species vocalizations and background environmental noise. This difference in recording conditions creates a domain shift, complicating the ability of models trained on focal data to generalize to soundscape recordings in PAM. In practice, we require models that can perform well across diverse and potentially unseen environments.
@@
Domain Generalization (DG) [8] aims to develop models that learn robust features that are domain-invariant, i.e. capable of generalizing to new unseen domains without prior knowledge or access to target domain data during training.
SupCon offers a promising approach for DG. In SupCon, the objective is to learn an embedding space where same-class examples are pulled together and different-class examples are pushed apart. This clustering can promote domain-invariance when sufficient domain diversity is present in the dataset, allowing the model to focus on features that are domain-invariant. In contrast, its self-supervised counterpart SimCLR [9] lacks this explicit mechanism for domain-invariance, as it relies solely on augmentations to create positive pairs. Without label information, SimCLR requires carefully designed augmentations that account for domain shift [10].
Despite its effectiveness, SupCon is computationally expensive due to the need for pairwise similarity calculations between all examples. To address this, we introduce ProtoCLR (Prototypical Contrastive Learning of Representations), a more efficient variant. By analyzing SupCon’s gradient and drawing inspiration from the generalization capabilities of prototypical networks [11] in few-shot learning, ProtoCLR replaces pairwise comparisons with a prototypical contrastive loss that compares examples to class prototypes, retaining the original objective while significantly reducing computational complexity.
@@ II Related Work
DG [8] has emerged as a critical approach to tackle domain shift, where the test data distribution differs from the training data. It aims to learn robust, domain-invariant representations using only source domain data, without requiring access to target domain data during training [8]. DG methods typically focus on learning domain-invariant representations, using techniques like domain alignment [17], meta-learning [18], and data augmentation [19]. These approaches help the model learn features that remain consistent across varying domains. In bioacoustics, DG is especially important due to the difficulty in collecting annotated soundscape recordings compared to focal data [12, 16].
Invariant learning has gained attention, where models are trained to learn features that remain invariant across different variations in data, such as augmented versions in self-supervised learning [9, 20, 21, 22] or same-class examples in supervised learning [5]. This approach has proven effective for learning robust features in bird sound classification [6].
III Method
III-A Supervised Contrastive Loss (SupCon)
@@
III-B Prototypical Contrastive Loss (ProtoCLR)
Motivated by the gradient of SupCon and drawing inspiration from prototypical networks [11], we propose ProtoCLR (Prototypical Contrastive Learning of Representations), which introduces class-level centroids into contrastive learning. The centroid for each class $y$ in the batch is computed as ${c}_{y}=\frac{\displaystyle 1}{\displaystyle|C(y)|}\sum_{i\in C(y)}{z}_{i}$, where $C(y)$ is the set of indices of examples with label $y$ and $|C(y)|$ is its size. We define the ProtoCLR loss as:
$\mathcal{L}^{\text{ProtoCLR}}=\sum_{i\in I}\frac{-1}{|P(i)|}\log\frac{\exp\left({z}_{i}\cdot{c}_{y_{i}}/\tau\right)}{\sum_{y\in Y}\exp\left({z}_{i}\cdot{c}_{y}/\tau\right)},$ (3)
@@ III-B Prototypical Contrastive Loss (ProtoCLR)
The positive term remains the same as in the gradient of SupCon, pulling the embeddings $\mathbf{z}_{i}$ towards the centroids of their respective classes. The difference is the negative term $-\frac{\displaystyle 1}{\displaystyle\tau}\frac{\sum_{y\in Y}S_{iy}{c}_{y}}{\sum_{y\in Y}S_{iy}}$: in ProtoCLR, the embeddings are pushed away from the weighted average of the centroids as opposed to the individual embeddings in SupCon.
III-C ProtoCLR vs SupCon
@@
III-C1 Complexity
SupCon has a computational cost of $\mathcal{O}(N^{2})$ due to computing dot products between all pairs of examples in a batch of size $N$, independent of the number of classes $C$. In contrast, ProtoCLR reduces this to $\mathcal{O}(N\times C)$ by computing dot products with class prototypes. Since $C$ is usually smaller than $N$, ProtoCLR is more efficient, particularly for large batches.
III-C2 Variance
@@
III-C3 Near Convergence Equivalence
Near convergence, both SupCon and ProtoCLR promote intra-class compactness with embeddings clustering tightly around class centroids: ${z}_{i}\approx{c}_{y_{i}}\quad\text{for all }i\in I$. In SupCon, the negative can be rewritten as:
@@ IV Experiments
TABLE I: Overview of BIRB datasets.
Model PER NES UHH HSN SSW SNE Mean
Random Guessing 0.75 1.12 3.70 5.26 1.04 1.78 2.22
@@
BirdAVES-biox-base 7.41$\pm$1.0 26.4$\pm$2.3 13.2$\pm$3.1 9.84$\pm$3.5 8.74$\pm$0.6 14.1$\pm$3.1 13.2
BirdAVES-bioxn-large 7.59$\pm$0.8 27.2$\pm$3.6 13.7$\pm$2.9 12.5$\pm$3.6 10.0$\pm$1.4 14.5$\pm$3.2 14.2
BioLingual 6.21$\pm$1.1 37.5$\pm$2.9 17.8$\pm$3.5 17.6$\pm$5.1 22.5$\pm$4.0 26.4$\pm$3.4 21.3
Perch 9.10$\pm$5.3 42.4$\pm$4.9 19.8$\pm$5.0 26.7$\pm$9.8 22.3$\pm$3.3 29.1$\pm$5.9 24.9
@@
CE 10.6$\pm$1.2 35.0$\pm$4.1 18.0$\pm$3.3 19.5$\pm$5.3 13.3$\pm$1.5 22.9$\pm$5.1 19.8
SimCLR 8.61$\pm$1.1 30.6$\pm$3.4 13.8$\pm$2.7 15.3$\pm$5.0 9.11$\pm$0.9 17.1$\pm$2.7 15.7
SupCon 7.87$\pm$0.9 34.5$\pm$2.5 15.6$\pm$3.4 18.5$\pm$4.9 12.1$\pm$1.5 19.6$\pm$3.7 18.0
ProtoCLR 8.30$\pm$1.1 36.5$\pm$3.4 15.2$\pm$3.0 19.8$\pm$5.6 11.7$\pm$1.6 20.3$\pm$3.3 18.6
@@
BirdAVES-biox-base 11.6$\pm$0.8 39.7$\pm$1.8 22.5$\pm$2.4 22.1$\pm$3.3 16.1$\pm$1.7 28.3$\pm$2.3 23.3
BirdAVES-bioxn-large 15.0$\pm$0.9 42.6$\pm$2.7 23.7$\pm$3.8 28.4$\pm$2.4 18.3$\pm$2.3 27.3$\pm$2.3 25.8
BioLingual 13.6$\pm$1.3 65.2$\pm$1.4 31.0$\pm$2.9 34.3$\pm$3.5 43.9$\pm$0.9 49.9$\pm$2.3 39.6
Perch 21.2$\pm$1.2 71.7$\pm$1.5 39.5$\pm$3.0 52.5$\pm$5.9 48.0$\pm$1.9 59.7$\pm$1.8 48.7
@@
CE 22.0$\pm$1.0 64.6$\pm$3.1 34.9$\pm$2.4 46.6$\pm$4.5 30.6$\pm$1.1 46.6$\pm$2.1 40.8
SimCLR 17.4$\pm$0.9 52.1$\pm$2.3 23.4$\pm$2.4 27.2$\pm$3.6 19.7$\pm$1.4 30.5$\pm$2.2 28.4
SupCon 16.1$\pm$1.0 60.3$\pm$2.0 29.7$\pm$2.5 35.7$\pm$5.6 25.9$\pm$1.2 39.4$\pm$2.7 34.5
ProtoCLR 16.5$\pm$1.1 61.3$\pm$2.3 30.0$\pm$2.4 38.4$\pm$4.4 26.0$\pm$1.3 39.5$\pm$2.3 35.3
TABLE II: Top-1 accuracy for one-shot and five-shot classification. All reported results are the average of ten runs. For our models, the results are further averaged over three checkpoints.
IV-A Few-Shot Classification Benchmark
The original task of BIRB [12] is information retrieval in bioacoustics, focused on retrieving bird vocalizations from passively recorded datasets using focal recordings for training. Downstream datasets may contain long audio recordings where events of interest are sparse. Individual bird events are detected using peak detection [23], which identifies frames with high energy to used for the task. We build on our previous work [6], which shows that using a pretrained classifier on AudioSet [24] is a good proxy for selecting windows with the highest bird class activation.
Following the MetaAudio framework, we focus on a multi-class classification task to define a few-shot evaluation for assessing the generalization capabilities of pre-trained models [25]. BIRB provides a large training set, XC (Xeno-Canto), consisting of focal recordings from nearly 10,000 species, while the validation and test datasets contain soundscape recordings. The benchmark details are provided in Table I.
@@ IV-A Few-Shot Classification Benchmark
To make predictions, we employ SimpleShot [27] by applying mean subtraction and L2-normalization to both centroids and test feature vectors. We then select the class whose centroid is closest to the test vector as the prediction. We repeat each few-shot experiment 10 times with different random seeds and report the mean and standard deviation accuracy in Table II.
IV-B Reference Systems
@@
To compare supervised contrastive approaches SupCon and ProtoCLR for addressing domain generalization, we train reference systems using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: BirdAVES [28], a transformer model based on the speech model HuBERT, trained in a self-supervised manner based on the AVES framework (explained in Section LABEL:motiv). We evaluate two versions of BirdAVES, biox-base, a 12-layer transformer with 768 units, trained on Xeno-Canto [xenocanto] and animal sounds from AudioSet [24] and VGGSound [vggsound], and bioxn-large, a 24-layer transformer with 1024 units, trained on the same datasets plus iNaturalist  [^2] https://github.com/gvanhorn38/iNatSounds. BirdAVES outputs one embedding vector per time frame, which we average across the time dimension to produce a single vector representing the entire audio clip; the encoder of BioLingual [29], an HT-SAT transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including Xeno-Canto, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [12], an EfficientNet-B1 [30] trained on Xeno-Canto for species classification, as well as taxonomic ranks genus, family, and order. We evaluate only models trained specifically on bird sounds, as general-purpose audio models have been shown to perform poorly in this domain [25]. Similarly, Hamer et al.[12] found that general-purpose models can, in some cases, perform worse than simply averaging mel spectrogram features without any learning.
IV-C Pre-training Details
@@
We follow the same preprocessing as Moummad et al.[6]. We train all models with CvT-13 [31], an efficient 2D transformer architecture with 20M parameters, on XC for 100 epochs using the AdamW optimizer with a batch size of 256, with a weight decay of $1\times 10^{-6}$. Following Moummad et al.[6], we apply the augmentations found to be effective for bird sound representations: circular time shift [32], SpecAugment [33], and spectrogram mixing [34]. These models are trained with a projector of dimension 128. For the CE loss, we only apply circular time shift and SpecAugment as augmentations, excluding Spectrogram Mixing, as it prevented the model from converging. The learning rate for CE and ProtoCLR is set to $5\times 10^{-4}$, while for SupCon and SimCLR, we use a learning rate of $1\times 10^{-4}$. We tune hyperparameters by monitoring k-NN accuracy on the POW dataset, which is split randomly into a training and a validation subset.
IV-D Results and Discussion
Table II presents the performance of different models on one-shot and five-shot bird sound classification tasks. ProtoCLR slightly outperforms SupCon both in one-shot and five-shot classification. Additionally, ProtoCLR is more computationally efficient; for one training epoch with a batch size of 256, SupCon computes 80.4B MACs, while ProtoCLR computes only 28.3B. On the other hand, CE outperforms ProtoCLR, when trained on 100 epochs. This could be due to the fact that contrastive methods require longer training iterations for optimal convergence than cross-entropy [5], which should be investigated.

### v6

@@ Abstract
Passive acoustic monitoring (PAM) is crucial for bioacoustic research, enabling non-invasive species tracking and biodiversity monitoring. Citizen science platforms provide large annotated datasets from focal recordings, where the target species is intentionally recorded. However, PAM requires monitoring in passive soundscapes, creating a domain shift between focal and passive recordings, challenging deep learning models trained on focal recordings. To address domain generalization, we leverage supervised contrastive learning by enforcing domain invariance across same-class examples from different domains. Additionally, we propose ProtoCLR, an alternative to SupCon loss which reduces the computational complexity by comparing examples to class prototypes instead of pairwise comparisons. We conduct few-shot classification based on BIRB, a large-scale bird sound benchmark to assess pre-trained bioacoustic models. Our findings suggest that ProtoCLR is a better alternative to SupCon. [^1] Models and code: https://github.com/ilyassmoummad/ProtoCLR
keywords:
contrastive learning, prototypical learning, domain generalization, few-shot learning, bioacoustics
1 Introduction
@@
However, a challenge arises from the difference between data collected on platforms like XC and PAM recordings. In citizen science platforms, recordings are typically focal, where the recorder is aimed directly at the species of interest. In contrast, PAM systems passively capture sounds within natural soundscapes, leading to recordings containing a mix of species vocalizations and background environmental noise. This difference in recording conditions creates a domain shift, complicating the ability of models trained on focal data to generalize to soundscape recordings in PAM.
@@
Domain Generalization (DG) [8] aims to develop models that learn robust features that are domain-invariant, i.e. capable of generalizing to new unseen domains without prior knowledge or access to target domain data during training. SupCon offers a promising approach for DG. In SupCon, the objective is to learn an embedding space where same-class examples are pulled together and different-class examples are pushed apart. This clustering can promote domain-invariance when sufficient domain diversity is present in the dataset, allowing the model to focus on features that are domain-invariant. In contrast, its self-supervised counterpart SimCLR [9] lacks this explicit mechanism for domain-invariance, as it relies solely on augmentations to create positive pairs. Without label information, SimCLR requires carefully designed augmentations that account for domain shift [10].
Despite its effectiveness, SupCon is computationally expensive due to the need for pairwise similarity calculations between all examples. To address this, in this paper we introduce ProtoCLR (Prototypical Contrastive Learning of Representations), a more efficient variant. By analyzing SupCon’s gradient and drawing inspiration from the generalization capabilities of prototypical learning [11, 12] in few-shot learning, ProtoCLR replaces pairwise comparisons with a prototypical contrastive loss that compares examples to class prototypes (one prototype per class), retaining the original objective while significantly reducing computational complexity.
@@ 2 Related Work
DG [8] has emerged as a critical approach to tackle domain shift, where the test data distribution differs from the training data. It aims to learn robust, domain-invariant representations using only source domain data.DG methods typically focus on learning domain-invariant representations, using techniques like domain alignment [18], meta-learning [19], and data augmentation [20]. These approaches help the model learn features that remain consistent across varying domains. In bioacoustics, DG is especially important due to the difficulty in collecting annotated soundscape recordings compared to focal data [13, 17].
Invariant learning has gained attention, where models are trained to learn features that remain invariant across different variations in data, such as augmented versions in self-supervised learning [9, 21, 22, 23] or same-class examples in supervised learning [5]. Moummad et al. [6] showed that invariant learning effectively learns representations of bird sounds that perform well in few-shot learning tasks for recognizing new species not seen during training. However, their experiments focus on 5-way 1-shot classification tasks, that do not represent realistic scenarios.In our study, we examine test datasets featuring passively recorded sounds from a variable number of species, reflecting more realistic and unencountered training scenarios.
3 Method
In this section, we revisit the supervised contrastive loss SupCon and derive ProtoCLR, an alternative that integrates class-level prototypes into the supervised contrastive learning paradigm. We discuss the computational advantages of ProtoCLR and show that they optimize for a similar objective near convergence.
Figure 1: Prototypical contrastive pre-training of representations: Anchors are augmented and pulled towards the prototype of their class, while being pushed apart from prototypes of other classes.
3.1 Supervised Contrastive Loss (SupCon)
@@
3.2 Prototypical Contrastive Loss (ProtoCLR)
Motivated by the gradient of SupCon and drawing inspiration from prototypical learning [11, 12], we propose ProtoCLR (Prototypical Contrastive Learning of Representations), which introduces class-level centroids into contrastive learning. The centroid for each class $y$ in the batch is computed as ${c}_{y}=\frac{\displaystyle 1}{\displaystyle|C(y)|}\sum_{i\in C(y)}{z}_{i}$, where $C(y)$ is the set of indices of examples with label $y$ and $|C(y)|$ is its size. We define the ProtoCLR loss as:
$\mathcal{L}^{\text{ProtoCLR}}=\sum_{i\in I}\log\frac{\exp\left({z}_{i}\cdot{c}_{y_{i}}/\tau\right)}{\sum_{y\in Y}\exp\left({z}_{i}\cdot{c}_{y}/\tau\right)},$ (3)
@@ 3.2 Prototypical Contrastive Loss (ProtoCLR)
The positive term remains the same as in the gradient of SupCon, pulling the embeddings ${z}_{i}$ towards the centroids of their respective classes. The difference is the negative term $\frac{\displaystyle 1}{\displaystyle\tau}\frac{\sum_{y\in Y}S_{iy}{c}_{y}}{\sum_{y\in Y}S_{iy}}$: in ProtoCLR, the embeddings are pushed away from the weighted average of the centroids as opposed to the individual embeddings in SupCon.
3.3 ProtoCLR vs SupCon
@@
3.3.1 Complexity
SupCon has a computational cost of $\mathcal{O}(N^{2})$ due to computing dot products between all pairs of examples in a batch of size $N$, independent of the number of classes $C$. In contrast, ProtoCLR reduces this to $\mathcal{O}(N\times C)$ by computing dot products with class prototypes. Since $C$ is smaller or equal to $N$, ProtoCLR is more efficient.
3.3.2 Variance
@@
3.3.3 Near Convergence Equivalence
Near convergence, both SupCon and ProtoCLR promote intra-class compactness with embeddings clustering tightly around class centroids: ${z}_{i}\approx{c}_{y_{i}}\quad\text{for all }i\in I$. In SupCon, the negative term in eq. (2) can be rewritten as:
@@ 4 Experiments
Table 1: Overview of BIRB datasets.
Model Training Params (M) PER NES UHH HSN SSW SNE Mean
Random Guessing - 0.75 1.12 3.70 5.26 1.04 1.78 2.22
@@
BirdAVES-biox-base 95 7.41$\pm$1.0 26.4$\pm$2.3 13.2$\pm$3.1 9.84$\pm$3.5 8.74$\pm$0.6 14.1$\pm$3.1 13.2
BirdAVES-bioxn-large 316 7.59$\pm$0.8 27.2$\pm$3.6 13.7$\pm$2.9 12.5$\pm$3.6 10.0$\pm$1.4 14.5$\pm$3.2 14.2
BioLingual 153 6.21$\pm$1.1 37.5$\pm$2.9 17.8$\pm$3.5 17.6$\pm$5.1 22.5$\pm$4.0 26.4$\pm$3.4 21.3
Perch 80 9.10$\pm$5.3 42.4$\pm$4.9 19.8$\pm$5.0 26.7$\pm$9.8 22.3$\pm$3.3 29.1$\pm$5.9 24.9
@@
CE 23 9.55$\pm$1.5 41.3$\pm$3.6 19.7$\pm$4.7 25.2$\pm$5.7 17.8$\pm$1.4 31.5$\pm$5.4 24.2
SimCLR 19 7.85$\pm$1.1 31.2$\pm$2.4 14.9$\pm$2.9 19.0$\pm$3.8 10.6$\pm$1.1 24.0$\pm$4.1 17.9
SupCon 19 8.53$\pm$1.1 39.8$\pm$6.0 18.8$\pm$3.0 20.4$\pm$6.9 12.6$\pm$1.6 23.2$\pm$3.1 20.5
ProtoCLR 19 9.23$\pm$1.6 38.6$\pm$5.1 18.4$\pm$2.3 21.2$\pm$7.3 15.5$\pm$2.3 25.8$\pm$5.2 21.4
@@
BirdAVES-biox-base 95 11.6$\pm$0.8 39.7$\pm$1.8 22.5$\pm$2.4 22.1$\pm$3.3 16.1$\pm$1.7 28.3$\pm$2.3 23.3
BirdAVES-bioxn-large 316 15.0$\pm$0.9 42.6$\pm$2.7 23.7$\pm$3.8 28.4$\pm$2.4 18.3$\pm$2.3 27.3$\pm$2.3 25.8
BioLingual 153 13.6$\pm$1.3 65.2$\pm$1.4 31.0$\pm$2.9 34.3$\pm$3.5 43.9$\pm$0.9 49.9$\pm$2.3 39.6
Perch 80 21.2$\pm$1.2 71.7$\pm$1.5 39.5$\pm$3.0 52.5$\pm$5.9 48.0$\pm$1.9 59.7$\pm$1.8 48.7
@@
CE 23 21.4$\pm$1.3 69.2$\pm$1.8 35.6$\pm$3.4 48.2$\pm$5.5 39.9$\pm$1.1 57.5$\pm$2.3 45.3
SimCLR 19 15.4$\pm$1.0 54.0$\pm$1.8 23.0$\pm$2.3 32.8$\pm$4.0 22.0$\pm$1.2 40.7$\pm$2.4 31.3
SupCon 19 17.2$\pm$1.3 64.6$\pm$2.4 34.1$\pm$2.9 42.5$\pm$2.9 30.8$\pm$0.8 48.1$\pm$2.4 39.5
ProtoCLR 19 19.2$\pm$1.1 67.9$\pm$2.8 36.1$\pm$4.3 48.0$\pm$4.3 34.6$\pm$2.3 48.6$\pm$2.8 42.4
Table 2: Top-1 accuracy for one-shot and five-shot classification. All reported results are the average of ten runs.
4.1 Few-Shot Classification Benchmark
The original task of BIRB [13] is information retrieval in bioacoustics, focused on retrieving bird vocalizations from passively recorded datasets using focal recordings for training. Downstream datasets may contain long audio recordings where events of interest are sparse. Individual bird events are detected using peak detection [24], which identifies frames with high energy to used for the task. We build on previous work [6], which shows that using a pretrained classifier on AudioSet [25] is a good proxy for selecting windows with the highest bird class activation.
Following the MetaAudio framework, we focus on a multi-class classification task to define a few-shot evaluation for assessing the generalization capabilities of pre-trained models [26]. BIRB provides a large training set, XC (Xeno-Canto), consisting of focal recordings from nearly 10,000 species, while the validation and test datasets contain soundscape recordings. The benchmark details are provided in Table 1.
Despite the inherently multi-label nature of soundscape recordings, we adopt a mono-class assumption for simplicity and consistency with the methodology used in BIRB [13]. To mitigate the effects of label confusion, we sample few-shot examples 10 times using different seeds.
@@ 4.1 Few-Shot Classification Benchmark
To make predictions, we employ SimpleShot [28] by applying mean subtraction and L2-normalization to both centroids and test feature vectors. We then select the class whose centroid is closest to the test vector as the prediction. We repeat each few-shot experiment 10 times with different random seeds and report the mean and standard deviation accuracy in Table 2.
4.2 Reference Systems
@@
To compare supervised contrastive approaches SupCon and ProtoCLR for addressing domain generalization, we train a reference system using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: BirdAVES [29], a transformer model based on the speech model HuBERT, trained in a self-supervised manner based on the AVES framework [29]. We evaluate two versions of BirdAVES, biox-base, a 12-layer transformer with 768 units, trained on Xeno-Canto [3] and animal sounds from AudioSet [25] and VGGSound [30], and bioxn-large, a 24-layer transformer with 1024 units, trained on the same datasets plus iNaturalist  [^2] https://github.com/gvanhorn38/iNatSounds. BirdAVES outputs one embedding vector per time frame, which we average across the time dimension to produce a single vector representing the entire audio clip; the encoder of BioLingual [31], an HT-SAT transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including Xeno-Canto, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [13], a large EfficientNet [32] trained on Xeno-Canto for species classification, as well as taxonomic ranks genus, family, and order. We evaluate only models trained specifically on bird sounds, as general-purpose audio models have been shown to perform poorly in this domain [26]. Similarly, Hamer et al. [13] found that general-purpose models can, in some cases, perform worse than simply averaging mel spectrogram features without any learning.
4.3 Pre-training Details
@@
We follow the same preprocessing as Moummad et al. [6]. We train all models with CvT-13 [33], an efficient 2D transformer architecture with 20M parameters, on XC for 300 epochs using the AdamW optimizer with a batch size of 256, with a weight decay of $1\times 10^{-6}$. Following Moummad et al. [6], we apply the domain-agnostic data augmentations found to be effective for bird sound representations: circular time shift [34], SpecAugment [35], and spectrogram mixing [36].These models are trained with a projector of dimension 128. For the CE loss, we only apply circular time shift and SpecAugment as augmentations, excluding Spectrogram Mixing, as it prevented the model from converging. The learning rate for CE and ProtoCLR is set to $5\times 10^{-4}$, while for SupCon and SimCLR, we use a learning rate of $1\times 10^{-4}$. We tune hyperparameters by monitoring k-NN accuracy on the POW validation dataset, which is split randomly into a training and a validation subset.
4.4 Results and Discussion
Table 2 presents the performance of different models on 1-shot and 5-shot bird sound classification tasks. ProtoCLR outperforms SupCon in both 1-shot and 5-shot classification.